OpenStudy (tatianagarcia90300):

solve the system of equations by substitution. what is the solution for x 2x + y = 1 4x + 2y = -1 A. x = 0 B. x = 2 C. there is no x value as there is no solution D. x can be any value as there are an infinite number of solutions

OpenStudy (freckles):

you can try to solve by elimination and see what happens multiply the first equation by -2 and then add the equations together afterwards

OpenStudy (freckles):

what did you when you multiply the first equation by -2?

OpenStudy (tatianagarcia90300):

y is equal to -2

OpenStudy (freckles):

I think you should try again,

OpenStudy (freckles):

multiply the first equation by -2 so -2(2x+y=1) is -4x-2y=-2 now add the following equations: -4x-2y=-2 4x+2y=-1 =============

OpenStudy (freckles):

add the first column -4x+4x add the second column -2y+2y = add the third column -2+(-1)

OpenStudy (tatianagarcia90300):

x + y = -3

OpenStudy (freckles):

something+(-same something)=0

OpenStudy (tatianagarcia90300):

so no solution

OpenStudy (freckles):

-x+x=0 x+(-x)=0 yes because we have first column=0 second column=0 = third column =-3 0+0=-3 <--not a true equation

OpenStudy (tatianagarcia90300):

oh ok i get it thank you so much

OpenStudy (freckles):

np